OpenStudy (jiteshmeghwal9):

If \[\sqrt{\frac{x}{1-x}}+\sqrt{\frac{1-x}{x}}={2{1 \over 2}}\] then the value of 'x' are:-

OpenStudy (anonymous):

have you tried solving for x?

OpenStudy (anonymous):

or is this like a trick question.

OpenStudy (anonymous):

Put: \[\frac{x}{1-x} = y\] Does it help??

OpenStudy (jiteshmeghwal9):

\[Ans: \frac{9}{13} , \frac{4}{13}\]

OpenStudy (jiteshmeghwal9):

I tried it but it doesn't work @Romero

OpenStudy (anonymous):

you might want to try waterineyes idea.

OpenStudy (jiteshmeghwal9):

K!

OpenStudy (anonymous):

Put: \[y = \sqrt{\frac{x}{1-x}}\] So, it becomes: \[y + \frac{1}{y} = \frac{5}{2}\] \[y^2 + 1 = \frac{5}{2}y\] \[2y^2 - 5y + 2 = 0\] Calculate y from here by using Quadratic Formula..

OpenStudy (anonymous):

Yeah just set it up like a general quadratic equation

OpenStudy (anonymous):

\[1 = \frac{x}{1-x}\] \[1 - x = x\] \[x = \frac{1}{2}\]

OpenStudy (anonymous):

Are you sure the answers are that you given??

OpenStudy (jiteshmeghwal9):

2y^2-5y+2=0\[\implies {-(-5)\pm \sqrt{(-5)^2-4(2)(2)}\over 2(2)}\]\[\implies {5\pm \sqrt{25-16} \over 4}\]\[\implies{5\pm \sqrt{9}\over 4}\]\[\implies{5\pm3 \over 4}\]. So, \[\alpha={5+3\over 4}\implies{8 \over 4}=2\] & \[Beta={5-3 \over 4} \implies{2 \over 4}={1 \over 2}\]

OpenStudy (jiteshmeghwal9):

Ya! I am 100 & 1 % sure:)

OpenStudy (anonymous):

Yes you are right.. It can be solved factorizing: \[(2y-1)(y-2) = 0\] \[y = 2 \quad Or \quad y = \frac{1}{2}\]

OpenStudy (jiteshmeghwal9):

But wht about my answer:(

OpenStudy (jiteshmeghwal9):

@mathslover Plz help:)

OpenStudy (anonymous):

We are going absolutely right but I don't know why answer is not right.. The answers you gave that are right..

OpenStudy (jiteshmeghwal9):

I also don't know :( @ParthKohli can u help???

mathslover (mathslover):

ok wait

OpenStudy (jiteshmeghwal9):

k!

OpenStudy (anonymous):

@jiteshmeghwal9 , There is \(2 \frac{1}{6}\) in the Right Hand Side, \[y = \sqrt{\frac{x}{1-x}}\] \[y^2 + 1 = \frac{13}{2}y\] \[6y^2 - 13y + 6 = 0\] \[y = \frac{13 \pm \sqrt}{}\] Using Quadratic Formula: \[y = \frac{2}{3} \quad Or \quad \frac{3}{2}\] Now solve for x you will definitely get your answer..

OpenStudy (jiteshmeghwal9):

Can u give me total answer:)

OpenStudy (anonymous):

Yes why not..

mathslover (mathslover):

wait i did mistake @waterineyes did i ?

OpenStudy (anonymous):

Where??

OpenStudy (jiteshmeghwal9):

@waterineyes Plz fast:)

OpenStudy (anonymous):

@jiteshmeghwal9 . The original equation is this: \[\sqrt{\frac{x}{1-x}}+\sqrt{\frac{1-x}{x}}={2{1 \over 6}}\] If you don't agree then put the answer in this and you can check..

OpenStudy (jiteshmeghwal9):

i agree.

mathslover (mathslover):